OpenStudy (math_nerd):

Let a,b,c be positive integers such that ab+bc-ca=0 and a-c=523. Find b.

ganeshie8 (ganeshie8):

what have you tried ?

OpenStudy (anonymous):

so this is a diff eq problem?

OpenStudy (anonymous):

set c= 523 plug them all in and try to find the answer

OpenStudy (mathmate):

@hzer I read \(a-c=523\)

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

well ok that is ok but also if you plug all of them in it will not give you the right one don't you think so mathmate

hartnn (hartnn):

this could be a start : \(\Large \dfrac{1}{b}=\dfrac{1}{a}+\dfrac{1}{c}\)

hartnn (hartnn):

b= c(c+523)/(2c+523)

hartnn (hartnn):

for b to be a positive integer 2c+523 should divide c (c+523)

OpenStudy (mathmate):

@hartnn I solved for c in c(c+523)/(2c+523)\(\in\) N [but by brute force]: a=137026=2*131*523 b=68382=2*3^2*29*131 c=136503=3^2*29*523 and 1/b=1/a+1/c ab+bc-ca=0 ...and there could be other solutions. I wish I knew how to solve for c "mathematically".